# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v3

@@ Abstract
Large Language Models (LLMs) such as GPT-4 have demonstrated their ability to understand natural language and generate complex code snippets. This paper introduces a novel Large Language Model Evolutionary Algorithm (LLaMEA) framework, leveraging GPT models for the automated generation and refinement of algorithms. Given a set of criteria and a task definition (the search space), LLaMEA iteratively generates, mutates and selects algorithms based on performance metrics and feedback from runtime evaluations. This framework offers a unique approach to generating optimized algorithms without requiring extensive prior expertise. We show how this framework can be used to generate novel black-box metaheuristic optimization algorithms automatically. LLaMEA generates multiple algorithms that outperform state-of-the-art optimization algorithms (Covariance Matrix Adaptation Evolution Strategy and Differential Evolution) on the five dimensional black box optimization benchmark (BBOB). The algorithms also show competitive performance on the 10- and 20-dimensional instances of the test functions, although they have not seen such instances during the automated generation process. The results demonstrate the feasibility of the framework and identify future directions for automated generation and optimization of algorithms via LLMs.
@@ I Introduction
Although they often provide algorithm design spaces of millions of potential module combinations plus their hyperparameter search spaces, modular frameworks also need to be created by experts in the field, by carefully selecting the included modules and providing the full infrastructure for configuring and searching through such algorithm spaces. Moreover, they are limited to the design space created by the choices that have been made by the experts when selecting modules for inclusion.
@@
In this work, we propose to overcome such limitations by using Large Language Models (LLMs) within an evolutionary loop for automatically and iteratively evolving and optimizing the program code of such metaphor-based optimization algorithms for solving the optimization problem in Equation 1. To test the approach, we use a specific optimization benchmarking tool (IOHprofiler, consisting of the IOHexperimenter [16] module for systematically running algorithms on benchmark functions, and IOHanalyzer [17] for statistically analyzing the results of the runs) to automatically evaluate the quality of the generated optimization algorithm and to provide a corresponding feedback to the LLM. The specific contributions of this work are as follows:
  • We present the LLM-Evolutionary Algorithm (LLaMEA), an evolutionary algorithm that uses an LLM to automatically generate and optimize high-quality metaphor-based optimization algorithms for solving the optimization problem as stated in Equation 1.
@@
  • We couple LLaMEA with the benchmarking tool, IOHexperimenter [17], to automatically evaluate the quality of the generated optimization algorithm, for providing feedback to the LLM. IOHexperimenter is a well-established benchmarking framework that supports evaluating optimization heuristics for continuous optimization problems; see e.g. [18, 19, 20, 21, 22]. It allows for comparing a new optimization algorithm automatically against a well-known set of state-of-the-art algorithms on a wide set of benchmark functions in a statistically sound way. It is only this automated benchmarking approach that allows us to close the evolutionary loop that includes the LLM and to run it automatically.
  • We demonstrate that LLaMEA can generate new metaphor-based optimization algorithms that successfully compete with or improve upon state-of-the-art algorithms on a standard set of 24 benchmark test problems (with multiple instances of each problem), the so-called black-box optimization benchmark BBOB [23], for the class of continuous optimization problem as stated in Equation 1. In particular, we show that the generated algorithms are competitive with the CMA-ES with default parameters, a hyperparameter-optimized CMA-ES, and Differential Evolution.
@@
We also would like to emphasize that the goal of this research is not to generate a specific algorithm for combinatorial or constraint optimization problems, but rather to combine the use of modern, sound benchmarking techniques for measuring the performance of algorithms with the generative capabilities of LLaMEA for achieving reliable results of the runs for the continuous optimization task defined in Equation 1. Going beyond this research, future work will address other application domains, too.
@@ II Related Work
Code generation: The recently proposed FunSearch approach (searching in the function space) [33] closes the loop and combines the LLM with a systematic evaluation of the quality of the generated programs, resulting in new solution algorithms for combinatorial optimization problems (cap-set problem, bin packing). In this approach, a distributed island-based Evolutionary Algorithm is used for maintaining diversity and prompt construction based on the already generated programs. In the approach called Algorithm Evolution using Large Language Model (AEL) [34] and also the Evolution of Heuristics (EoH) [35], ideas from Evolutionary Algorithms are used explicitly to design novel optimization heuristics. AEL and EoH treat each algorithm as a solution candidate in a population, evolving them by asking the LLM to perform crossover and mutation operators on the heuristics generated by the LLM. These approaches have shown superior performance over traditional heuristic methods and domain-specific models in solving small instances of the Traveling Salesperson Problem (TSP). The AEL approach, however, is very specific to the TSP and cannot be applied to the continuous optimization problem as in Equation 1, which is the subject of our work. The EoH approach can be generalized to the continuous optimization problem, and thus be compared with our proposed approach. Note that the EoH approach was designed to generate and optimize small pieces of code (single functions) and the black-box optimization algorithms we aim to generate can be complex, relatively large and consisting of a complete class with variables and functions.
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite (see [30] and the supplemental material for an overview of the functions). The suite consists of $24$ noiseless functions, with an instance generation mechanism to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into five function groups: (i) separable functions, (ii) functions with low or moderate conditioning, (iii) high conditioning unimodal functions, (iv) multi-modal functions with strong global structure and (v) multi-modal functions with weak global structure. The BBOB suite is considered a best-practice in the field of metaheuristic algorithm design for the evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, namely the normalized Area Over the Convergence Curve (AOCC) (see Equation 2). This results in one number $y_{t}$ representing aggregated performance of the proposed algorithm over the complete benchmark suite, including multiple instances per function, and its standard deviation $\sigma_{t}$ over multiple runs of the algorithm.
@@ III-D Mutation, Selection and Feedback
  Feedback prompt template $F$     <Task prompt S>  <List of previously generated algorithm names with mean AOCC score>  <selected algorithm to refine (full code) and mean and std AOCC scores>  Either refine or redesign to improve the algorithm.
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean AOCC values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$. In our experiments we cover the inclusion of such details versus leaving them out.
@@
In the proposed framework we do not keep the entire run history in every iteration (including all codes $(a_{0},\ldots,a_{t})$), as this becomes more and more expensive as the list grows. However, the inclusion of a larger set of previous attempts by providing the previous solutions in code with their associated scores could in theory be beneficial. The LLM would be able to act as a kind of surrogate model in predicting the next solution. This would translate to evolutionary computation terms, as to keeping an archive of best solutions or the use of machine learning as surrogate models in surrogate assisted optimization [40, 41].
@@ IV Experimental Setup
To validate the proposed evolution framework for generating and optimizing metaheuristics, we designed a set of experiments to compare different LLMs, different ES strategies and different levels of detail in the provided evaluation feedback. In addition, the best generated algorithms are analyzed and compared to several state-of-the-art baselines on the BBOB suite using additional instances and additional dimension settings.
@@ IV-C Performance Metrics
To evaluate the generated algorithms effectively over a complete set of benchmark functions we use a so-called anytime performance measure, meaning that it quantifies performance of the optimization algorithm over the complete budget, instead of only looking at the final objective function value. For this we use the normalized Area Over the Convergence Curve (AOCC), as introduced in [45]. The AOCC is given in Equation 2.
@@
We aggregate the AOCC scores of all 24 BBOB benchmark functions $f_{1},\ldots,f_{24}$ by taking the mean over functions and their instances, i.e., for an algorithm $a$,
$\textit{AOCC}(a)=\frac{1}{3\cdot 24}\sum_{i=1}^{24}\sum_{j=1}^{3}\textit{AOCC}(\mathbf{y}_{a,f_{ij}})\;.$ (3)
The final mean AOCC over $k=5$ independent runs of algorithm $a$ over all BBOB functions is given as feedback to the LLM in the next step, and is used as best-so-far solution in case an improvement was found. In other words, the quality measure $f(a)$ used in Algorithm 1 is defined as
@@ V Results and Discussion
In Figure 2, the median best-so-far algorithm evaluation score (AOCC) per configuration (LLM and feedback strategy) is shown.
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and strategy, including the state-of-the-art baseline EoH algorithm. Shaded areas denote the variance of the best-so-far. w/Details denotes that we use a feedback mechanism that provides not just the plain average AOCC (and its standard deviation) but also the average AOCC and standard deviation per BBOB function group as feedback to the LLM.
@@
From Figure 2 we conclude that different strategies yield a range of different results, depending on the model used. For example, the $(1+1)$-selection strategy seems to be beneficial for GPT-4, but is having a deteriorating effect when using GPT-4o. Overall GPT-4 seems to be better suited for the task overall, and GPT-3.5 is clearly a less favorable option. We also compare to the EoH algorithm with GPT-4 as our baseline. To use EoH, we extended it to work in tandem with IOHexperimenter in the same way as LLaMEA. The $(1+1)$-GPT-4-ES shows better convergence (area under the AOCC curve) than the EoH approach and also results in an overall better algorithm (AOCC) in the end. The EoH approach follows as the second best approach, EoH is run $5$ times with a population size of $5$ and default hyperparameters. EoH performs a bit more stable (in terms of variation over different runs), due to the larger population size. However, it suffers from tries where the generated code is not executable since it has no self-debugging capabilities (unlike LLaMEA). Another limitation of EoH is that it can only deal with single functions, while in our proposed approach complete Python classes are generated, allowing for more complex interactions.
@@ V-A Novelty and diversity
In Figure 3, a word-cloud is shown with all the sub-strings of generated algorithm names and their occurrence frequency visualized. Note that the word-cloud serves purely as an intuitive visual representation to quickly get an overview of the different words the algorithm names contain. All algorithm names and codes are available in our open-source repository [39]. The generated algorithm names are in Python Camel-case style, such that it is easy to split the algorithm names into individual parts. In some cases, the algorithm name is just an abbreviation, and these abbreviations are kept as words in this analysis. The most used parts in algorithm names are rather generic, such as “evolution”, “adaptive” and “dynamic”. Some of the parts directly refer to existing algorithms, such as “harmony” and “firework”, and other strings refer to existing strategies, such as “gradient”, “local”, “elite” etc. In general when observing the different generated solutions, we see interesting and novel combinations of existing techniques, such as ”surrogate assisted differential evolution combined with covariance matrix adaptation evolution strategies” and ”dynamic firework optimizer with enhanced local search”. A full list of generated algorithms and their names is provided in our Zenodo repository [39]. The algorithm names are generated automatically by the LLM that is used by LLaMEA. We have performed manual checks and can confirm that the algorithm names are almost always in line with the actual code. This means that the LLM indeed creates a descriptive name for the generated algorithm, which is based on the key algorithmic components used in the algorithm.

### v4

@@ Abstract
Large Language Models (LLMs) such as GPT-4 have demonstrated their ability to understand natural language and generate complex code snippets. This paper introduces a novel Large Language Model Evolutionary Algorithm (LLaMEA) framework, leveraging GPT models for the automated generation and refinement of algorithms. Given a set of criteria and a task definition (the search space), LLaMEA iteratively generates, mutates and selects algorithms based on performance metrics and feedback from runtime evaluations. This framework offers a unique approach to generating optimized algorithms without requiring extensive prior expertise. We show how this framework can be used to generate novel black-box metaheuristic optimization algorithms for box-constrained, continuous optimization problems automatically. LLaMEA generates multiple algorithms that outperform state-of-the-art optimization algorithms (Covariance Matrix Adaptation Evolution Strategy and Differential Evolution) on the 5-dimensional black box optimization benchmark (BBOB). The algorithms also show competitive performance on the 10- and 20-dimensional instances of the test functions, although they have not seen such instances during the automated generation process. The results demonstrate the feasibility of the framework and identify future directions for automated generation and optimization of algorithms via LLMs.
@@ I Introduction
Although they often provide algorithm design spaces of millions of potential module combinations plus their hyper-parameter search spaces, modular frameworks also need to be created by experts in the field, by carefully selecting the included modules and providing the full infrastructure for configuring and searching through such algorithm spaces. Moreover, the modular approaches are limited to the design space created by the choices that have been made by the experts when selecting modules for inclusion.
@@
In this work, we propose to overcome such limitations by using Large Language Models (LLMs) within an evolutionary loop for automatically and iteratively evolving and optimizing the program code of such metaheuristic optimization algorithms for solving the optimization problem in Equation (1). To evaluate the generated algorithms, we use a specific optimization benchmarking tool (IOHprofiler, consisting of the IOHexperimenter [16] module for systematically running algorithms on benchmark functions, and IOHanalyzer [17] for statistically analyzing the results of the runs) to automatically evaluate the quality of the generated optimization algorithms and to provide a corresponding feedback to the LLM. The specific contributions of this work are as follows:
  • We present the LLM-Evolutionary Algorithm (LLaMEA), an evolutionary algorithm that uses an LLM to automatically generate and optimize high-quality metaheuristic optimization algorithms for solving the optimization problem as stated in Equation (1). LLaMEA combines in-context learning, a precise task prompt (including a sample algorithm, code interface requirements, and mutation instruction), error handling and two different selection strategies in a novel way. To the best of our knowledge, this is the first time that new metaheuristics for continuous optimization problems are successfully generated and optimized by LLMs to reach and exceed state-of-the-art optimization algorithm performance levels. In particular, we show that the generated algorithms are competitive with three baseline algorithms, namely (i) the CMA-ES with default parameters, (ii) a hyper-parameter-optimized CMA-ES, and (iii) Differential Evolution.
@@
  • We couple LLaMEA with the benchmarking tool, IOHexperimenter [17], to automatically evaluate the quality of the generated optimization algorithm, for providing feedback to the LLM. IOHexperimenter is a well-established benchmarking framework that supports evaluating optimization heuristics for continuous optimization problems; see e.g. [18, 19, 20, 21, 22]. It allows for comparing a new optimization algorithm automatically against a well-known set of state-of-the-art algorithms on a wide set of benchmark functions in a statistically sound way. It is only this automated benchmarking approach that allows us to close the evolutionary loop that includes the LLM and to run it automatically. We specifically use a standard set of 24 benchmark test problems (with multiple instances of each problem), the so-called black-box optimization benchmark BBOB [23], for representing the class of continuous optimization problem as stated in Equation (1).
  • To evaluate the quality of a newly generated metaheuristic across a set of benchmark functions by a single, aggregated performance measure, we use the Area Over the Convergence Curve ($AOCC$) measure [24]. This enables the LLM to generate metaheuristics that perform well across a set of problem instances, rather than on a single problem, only.
@@
We also would like to emphasize that the goal of this research is not to generate a specific algorithm for combinatorial or constraint optimization problems, but rather to combine the use of modern, sound benchmarking techniques for measuring the performance of algorithms with the generative capabilities of LLaMEA for achieving reliable results for the continuous optimization task defined in Equation (1) and for outperforming state-of-the-art metaheuristics on such problems.
@@ II Related Work
Code generation: The recently proposed FunSearch approach (searching in the function space) [34] closes the loop and combines the LLM with a systematic evaluation of the quality of the generated programs, resulting in new solution algorithms for combinatorial optimization problems (cap-set problem, bin packing) after 2.5 million calls of the LLM. In this approach, a distributed island-based Evolutionary Algorithm is used for maintaining diversity and prompt construction based on the already generated programs. In the approach called Algorithm Evolution using Large Language Model (AEL) [35] and also the Evolution of Heuristics (EoH) [36], ideas from Evolutionary Algorithms are used explicitly to design novel optimization heuristics. AEL and EoH treat each algorithm as a solution candidate in a population, evolving them by asking the LLM to perform crossover and mutation operators on the heuristics generated by the LLM. These approaches have shown superior performance over traditional heuristic methods and domain-specific models in solving small instances of the Traveling Salesperson Problem (TSP). The AEL approach, however, is very specific to the TSP and cannot be applied (trivially) to the continuous optimization problem as in Equation (1), which is the subject of our work. The EoH approach can be generalized to the continuous optimization problem with some small modifications, and thus be compared with our proposed approach. Note that the EoH approach was designed to generate and optimize small pieces of code (single functions) and the black-box optimization algorithms we aim to generate can be complex, relatively large and consisting of a complete class with variables and functions.
@@ III-C Evaluation
The generated algorithm is evaluated on the Black-Box Optimization Benchmark (BBOB) suite (see [31] and the supplemental material for an overview of the functions). The suite consists of $24$ noiseless functions, with an instance generation mechanism to generate a diverse set of different optimization landscapes that share particular function characteristics. The suite is divided into five function groups: (i) separable functions, (ii) functions with low or moderate conditioning, (iii) high conditioning unimodal functions, (iv) multi-modal functions with strong global structure and (v) multi-modal functions with weak global structure. The BBOB suite is considered a best-practice in the field of metaheuristic algorithm design for the evaluation of newly proposed algorithms. For the evaluation feedback provided to the LLM, we summarize the performance on the whole BBOB suite by taking an average any-time performance metric, namely the normalized Area Over the Convergence Curve (AOCC) (see Equation (2)). This results in one number $y_{t}$ representing aggregated performance of the proposed algorithm over the complete benchmark suite, including multiple instances per function, and its standard deviation $\sigma_{t}$ over multiple runs of the algorithm.
@@ III-D Mutation, Selection and Feedback
  Feedback prompt template $F$     <Task prompt S>  <List of previously generated algorithm names with mean \$AOCC\$ score>  <selected algorithm to refine (full code) and mean and std \$AOCC\$ scores>  Either refine or redesign to improve the algorithm.
@@
Instead of providing an overall score and standard deviation, we can provide the LLM with more performance details of the generated algorithms. In our case, we can provide additional metrics per BBOB function group, potentially giving the LLM information on what kind of functions the solution works well and on which ones it does not work well. This results in mean $AOCC$ values and standard deviations for each of the five function groups, i.e., $(y_{t,1},\ldots,y_{t,5}),(\sigma_{t,1},\ldots,\sigma_{t,5})$.
In our experiments we cover the inclusion of such details versus leaving them out. However, since adding such details never showed an increase in performance, we have left these results in the supplemental material to keep our main results concise and clear.
@@
In the proposed framework we do not keep the entire run history in every iteration (including all codes $(a_{0},\ldots,a_{t})$), as this becomes more and more expensive as the list grows. However, the inclusion of a larger set of previous attempts by providing the previous solutions in code with their associated scores could in theory be beneficial. The LLM would be able to do multi-shot in-context learning, and potentially learn from previously generated solutions. This would translate to evolutionary computation terms, as to keeping an archive of best solutions or the use of predictive machine learning models in machine learning assisted optimization [41, 42].
@@ IV Experimental Setup
To validate the proposed evolution framework for generating and optimizing metaheuristics, we designed a set of experiments to compare three different LLMs and the two different selection strategies for each of them. We then compare the best of these six combinations of LLM and selection strategy with the state of the art EoH algorithm (using the best-performing LLM) and a random search baseline (also using the best-performing LLM). In addition, the best generated algorithms are analyzed and compared to several state-of-the-art baselines on the BBOB suite using additional instances and additional dimension settings.
@@ IV-C Performance Metrics
To evaluate the generated algorithms effectively over a complete set of benchmark functions we use a so-called anytime performance measure, meaning that it quantifies performance of the optimization algorithm over the complete budget, instead of only looking at the final objective function value. For this we use the normalized Area Over the Convergence Curve ($AOCC$), as introduced in [46]. The $AOCC$ is given in Equation (2).
@@
Following common practice, we aggregate the $AOCC$ scores of all 24 BBOB benchmark functions $f_{1},\ldots,f_{24}$ by taking the mean over functions and their instances, i.e., for an algorithm $a$,
$\textit{AOCC}(a)=\frac{1}{3\cdot 24}\sum_{i=1}^{24}\sum_{j=1}^{3}\textit{AOCC}(\mathbf{y}_{a,f_{ij}})\;,$ (3)
where $f_{ij}$ is the $j$-th instance of function $i$. The final mean $AOCC$ over $k=5$ independent runs of algorithm $a$ over all BBOB functions is given as feedback to the LLM in the next step, and is used as best-so-far solution in case an improvement was found. In other words, the quality measure $f(a)$ used in Algorithm 1 is defined as
@@
IV-D Baselines
To assess the performance of the proposed framework, we compare it with two baselines, namely the state-of-the-art approach Evolution of Heuristics (EoH), and random search (RS), also evaluated on the BBOB benchmark function suite using the same performance measure based on $AOCC$, given in Equation (4). For EoH to work on our use-case, we modified the approach slightly, making as little changes as possible. We keep the same prompt as we used for the proposed approach, with the exception that it should generate a ”small description” instead of a name and that it should generate ”a single function with a specific name” (both required by EoH to work). We then linked the evaluation function to IOHexperimenter in the exact same way as for the proposed approach. All EoH hyper-parameters are kept at their default values. Since EoH is minimizing by default we multiply the evaluated fitness by $-1$ before returning it to EoH. The random search baseline is just prompting the LLM for $100$ times using the same starting prompt. Both baselines use the GPT-4 LLM as this was the best-performing choice.
@@ V Results and Discussion
In Figure 2, the mean best-so-far algorithm evaluation score ($AOCC$ according to Equation (4)) per configuration (LLM and strategy) is shown. Since we show only best-so-far values, all curves are monotonically increasing, also when (1,1)-selection is used. When execution errors occur, which happens in $18.7$% ($843$ out of 4 50045004\,5004 500) of all runs, the corresponding $AOCC$ value of zero is not plotted, as this would make the graphs unreadable.
Figure 2: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for each LLM and selection strategy. Shaded areas denote the standard deviation of the best-so-far. Please note that the difference in initial performance already results from the fact that, although the starting prompt $S$ is identical for all LLMs, the performance value shown here is the mean $AOCC$ value of the first algorithm $a_{1}$ generated by each LLM in line 4 of Algorithm 1. Notice that only best-so-far values are plotted, also for (1,1)-strategies, and infeasible algorithm results are also not plotted (as they would have an $AOCC=0$ value as mentioned in Algorithm 1).
Figure 3: Mean convergence curves (best-so-far algorithm scores) over the $5$ different runs for the best strategy LLaMEA-variant, i.e., LLaMEA-(1+1) GPT-4 (same curve as in Fig. 2), including the state-of-the-art baseline EoH algorithm (red) and the random search (RS) baseline (black). Shaded areas denote the standard deviation of the best-so-far over $5$ runs. Additional remarks as provided in the caption of Fig. 2 apply here, too.
@@
From Figure 2 we conclude that different strategies yield a range of different results, depending on the model used. For example, the $(1+1)$-selection strategy seems to be beneficial for GPT-4, but is having a deteriorating effect when using GPT-4o. Overall GPT-4 seems to be better suited for the task overall, and GPT-3.5 is clearly a less favorable option. In Figure 3, we compare our approach with the two baseline approaches, namely the EoH algorithm and random search (each using GPT-4, too). To use EoH, we extended it to work in tandem with IOHexperimenter in the same way as LLaMEA. The LLaMEA-$(1+1)$ GPT-4 shows better convergence (area under the $AOCC$ curve) than the EoH approach and also results in an overall better algorithm in the end. Both EoH and the proposed approach show a clear improvement over just randomly sampling the LLM, indicating the advantages of using an evolutionary search framework. The EoH approach follows as the second best approach, where EoH is also run $5$ times with a population size of $5$ and default hyper-parameters. EoH performs a bit more stable (in terms of variation over different runs), due to the larger population size. However, it suffers from tries where the generated code is not executable since it has no self-debugging capabilities (unlike LLaMEA). Another limitation of EoH is that it can only deal with single functions, while in our proposed approach complete Python classes are generated, allowing for more complex interactions.
@@ V-A Novelty and diversity
In Figure 4, a word-cloud is shown with all the sub-strings of generated algorithm names and their occurrence frequency visualized. Note that the word-cloud serves purely as an intuitive visual representation to quickly get an overview of the different words the algorithm names contain. All algorithm names and codes are available in our open-source repository [40]. The generated algorithm names are in Python Camel-case style, such that it is easy to split the algorithm names into individual parts. In some cases, the algorithm name is just an abbreviation, and these abbreviations are kept as words in this analysis. The most used parts in algorithm names are rather generic, such as “evolution”, “adaptive” and “dynamic”. Some of the parts directly refer to existing algorithms, such as “harmony” and “firework”, and other strings refer to existing strategies, such as “gradient”, “local”, “elite” etc. In general when observing the different generated solutions, we see interesting and novel combinations of existing techniques, such as “surrogate assisted differential evolution combined with covariance matrix adaptation evolution strategies” and “dynamic firework optimizer with enhanced local search”. A full list of generated algorithms and their names is provided in our Zenodo repository [40]. The algorithm names are generated automatically by the LLM that is used by LLaMEA. We have performed manual checks and can confirm that the algorithm names are almost always in line with the actual code. This means that the LLM indeed creates a descriptive name for the generated algorithm, which is based on the key algorithmic components used in the algorithm.